### v1

# Multivalued (Wess-Zumino-Novikov) Functional in Fluid Mechanics

Abstract
We introduce hydrodynamics derived from the action of a perfect fluid deformed by the Novikov multivalued functional generated by combined actions of gauge transformations and spacetime diffeomorphisms. This deformation is structured analogously to the Wess-Zumino functional, commonly used as an effective description of field theories with the chiral anomaly. The deformation affects the continuity equation giving fluid spin degrees of freedom and is consistent with the kinematics of the chiral anomaly.
1. Introduction and the Main Result.
The equations of motion for a perfect fluid can be regarded as conservation laws associated with the group of spacetime diffeomorphisms. If no external forces act on the fluid the momentum and energy are conserved resulting in four dynamical equations expressed as a divergence-free condition of the canonical momentum-stress-energy tensor
$\partial_{\mu}\tensor{T}{{}^{\mu}_{\nu}}=0\,.$ (1)
These equations alone are sufficient for describing a homentropic and barotropic flows, special flows, when the only dynamical variables are components of the particle number 4-current $n^{\mu}=(n^{0},\,n^{0}\bm{v})$, where $n^{0}$ is the particle number density and $\bm{v}$ is the fluid velocity. In this case the continuity equation
$\partial_{\mu}n^{\mu}=0\,$ (2)
is not an independent condition. It follows from the conservation of momentum and energy described by (1).
A more general, baroclinic flow involves additional dynamical variables. In this case, more equations stem from symmetries other than spacetime diffeomorphisms are necessary. Among them is the gauge symmetry, which yields the continuity equation. In the usual perfect fluid the continuity equation means the particle number conservation (2).
The phenomenon known as the chiral current anomaly presents an obstacle to the conservation of particle number. The issue arises when the Noether current generated by the gauge symmetry, denoted as $I^{\mu}$, is not gauge-invariant; however, its divergence is. In this case, the particle number is not conserved $\partial_{\mu}n^{\mu}\neq 0$, while the equations of motion remain local and gauge-invariant. The chiral anomaly signifies that the flow entrains a reservoir capable supplying and swapping particles.
The chiral anomaly was initially identified as a kinematic property of quantum field theories involving chiral (or Weyl) fermions Current_algebra . A defining feature of the chiral anomaly is that the particle production rate, $\partial_{\mu}n^{\mu}$, is locally defined by the flow itself and is unaffected by changes in the spacetime metric. Therefore, the anomaly is largely insensitive to interaction and, when carried over to a liquid state it does not introduce additional spacetime scales beyond already accounted gradients of hydrodynamic fields. Being insensitive to a variation of metric, the chiral anomaly only impacts the continuity equation while leaving the form of the stress tensor unaffected.
In recent years, there has been growing confidence that the current anomalies are compatible with classical fluid dynamics. An incomplete list of references is volovik ; Erdmenger ; son2009hydrodynamics ; stephanov2012chiral ; jensen2013thermodynamics ; haehl2014effective ; stephanov_son2015 ; AMN ; abanov2022axial ; wiegmann2022chiral ; AW2 ; wiegmann2023Hamilton ; Abanov_Cappelli . A physical argument supporting this perspective is the existence of liquids composed of Weyl fermions. Such liquids are expected to retain the kinematics of Weyl fermions, including their anomalies. A notable example is the superfluid ${\rm He}^{3}$ volovik .
In the case, of a single component fluid, an ‘electric current’, which is the Noether current generated by the gauge symmetry would be a particle number current $n^{\mu}$ modified by an ’anomalous’ pseudo-vector field $h^{\mu}$
$I^{\mu}=n^{\mu}+\tfrac{k}{2}h^{\mu}\,,$ (3)
where $k$ represents the strength of the deformation. The vector field $h^{\mu}$ does not depend on the particle number and (i) is locally expressed in terms of the momentum of the flow, (ii) is metric independent, and (iii) its divergence is gauge-invariant. Then the gauge-invariant condition
$\partial_{\mu}I^{\mu}=0\,$ (4)
replaces the continuity equation (2).
In this paper, we seek a minimal deformation of the hydrodynamics of a perfect fluid with these properties. We remark that a room for deformations of the equations of fluid dynamics is limited, as they must maintain covariance under the action of the gauge group and the group of spacetime diffeomorphisms
$\mathcal{G}=U(1)\rtimes{\rm\ Diff}(\mathcal{M}^{4})\,.$ (5)
The Hamilton principle of hydrodynamics can serve as a concise criterion to ensure that the equations of motion maintain this property. The Hamilton principle asserts the existence of a Hamilton functional, whose invariance under the action of the symmetry group $\cal G$ yields the equations of fluid dynamics. Our strategy follows thus: we deform the Hamilton function, and subsequently derive the equations of motion.
The deformation of the Hamilton functional is carried out by the multivalued functional introduced by Novikov in 1981 Novikov . Soon after Novikov’s paper, it was recognized that a class of these functionals already appeared in the early work of Wess and Zumino Wess-Zumino , albeit in a coordinate form. Wess and Zumino constructed the functional whose variation replicates the l chiral anomaly effects in the linear response to external gauge field of Weyl fermions. The multivalued functional suitable for hydrodynamics stems from treating the fluid phase space as a manifold of the infinite-dimensional Lie group $\cal G$, in line with Arnold’s approach to hydrodynamics Arnold .
The anomaly is a topological phenomenon in the sense that it is metric-independent. For this reason, a spacetime covariant formulation of hydrodynamics developed by Lichnerowicz Lichnerowicz and Carter Carter provides a concise and efficient framework for analyzing conservation laws in fluid dynamics (see Gourgoulhon ; Markakis for a more recent review). In this approach, the hydrodynamics is expressed in terms of the particle number 4-current $n^{\mu}$ and its canonical conjugate, the kinetic fluid 4-momentum $p_{\mu}$, without reference to the spacetime metric. Consequently, the formulas appear identical for both relativistic and non-relativistic fluids. The metric enters through the covariant version of the ‘equation of state’ which relates $n^{\mu}$ and $p_{\mu}$. Our results are not sensitive to it.
For instance, the canonical stress tensor of a perfect fluid in (1) expressed in canonical variables is
$\tensor{T}{{}^{\mu}_{\nu}}=n^{\mu}p_{\nu}+\tensor{\delta}{{}^{\mu}_{\nu}}P\,,$ (6)
where $P$ is the fluid pressure.
It’s not difficult to identify the anomalous current in equation (3), given that the fluid phase space presents only one current distinct from the particle number current. This is the fluid helicity Erdmenger ; son2009hydrodynamics
$h^{\mu}=\epsilon^{\mu\nu\lambda\sigma}\pi_{\nu}\partial_{\lambda}\pi_{\sigma}\,,$ (7)
the 3-form constructed from of fluid canonical 4-momentum $\pi_{\mu}$. Compared to kinetic momentum $p_{\mu}$, the canonical momentum is not gauge invariant. At no external gauge field, such as an electromagnetic field, the two momenta are related by a gradient of a gauge phase
$\pi_{\mu}=p_{\mu}+\partial_{\mu}\psi\,.$ (8)
In a typical fluid, the gauge phase has no physical significance. However, with the chiral anomaly, the scalar $\psi$ takes on a physical meaning. It does not factor into the equation of motion, but enters the Hamilton functional (see Eq.(40) below).
We can express the Noether current (3) in terms a gauge-invariant kinetic helicity
$\Sigma^{\mu}:=\epsilon^{\mu\nu\lambda\sigma}p_{\nu}\partial_{\lambda}p_{\sigma}\,.$ (9)
as
$I^{\mu}=j^{\mu}+\tfrac{k}{2}\epsilon^{\mu\nu\lambda\sigma}\partial_{\nu}\psi\,\partial_{\lambda}\pi_{\sigma}\,,$ (10)
where $j^{\mu}$ is the gauge invariant part of the current
$j^{\mu}=n^{\mu}+\tfrac{k}{2}\Sigma^{\mu}\,.$ (11)
The non-gauge-invariant term in (10) represents the part of the ‘electric’ current flowing through the reservoir. As a curl, it drops out of the continuity equation (4), which, when expressed solely in gauge-invariant terms, becomes
$\partial_{\mu}j^{\mu}=0\,.$ (12)
The effects of the chiral anomaly are even more apparent when the fluid is charged and exposed to an electromagnetic field. While we do not explore these effects here, we list the equation of motion without providing the derivation. The gauge potential modifies the canonical momentum as $\pi_{\mu}=p_{\mu}+A_{\mu}$, thereby altering the kinetic helicity $\Sigma^{\mu}=\epsilon^{\mu\nu\lambda\sigma}p_{\nu}(\partial_{\lambda}p_{\sigma}+F_{\lambda\sigma})$ by the Maxwell field tensor $F_{\mu\nu}$. Consequently, the equations of motion (1,12) transform into
$\partial_{\mu}\tensor{T}{{}^{\mu}_{\nu}}=F_{\nu\mu}\,j^{\nu},\quad\partial_{\mu}j^{\mu}=\tfrac{k}{4}\,\,\epsilon^{\mu\nu\lambda\sigma}F_{\nu\mu}{}^{\star}F^{\mu\nu}\,,$ (13)
where ${}^{\star}F^{\mu\nu}=\tfrac{1}{2}\epsilon^{\mu\nu\lambda\sigma}F_{\lambda\sigma}$ is the dual tensor. The second relation, which replaces (12), is commonly known as chiral anomaly.
Our construction holds for any even spacetime dimensions, as the anomalous term in (3) remains as helicity. Expressed in terms of differential forms $\pi=\pi_{\mu}dx^{\mu}$ the anomalous term is given by
$\frac{k}{(d/2)!}\pi\wedge(d\pi)^{d/2-1}\,.$ (14)
In particular, this formula reproduces the known expression for the Noether current of (1+1)-chiral bosons
$I^{\mu}=n^{\mu}+k\epsilon^{\mu\nu}\pi_{\nu}\,.$ (15)
Before we can address the multivalued functional, it is necessary to outline the hydrodynamic setup.
2. Semiclassical Quantization of Fluid Helicity, Vortex Linking and Particle Production.
We begin with a remark on the normalization of the fluid helicity (7) and interpretation of the (yet to obtain) the deformed continuity equation (4).
While our discussion primarily focuses on classical fluid dynamics, a natural normalization arises in semiclassical fluids. We recall that in semiclassical fluids vorticity is localized in vortex lines (or loops in the absence of spatial boundaries), with the vortex circulation being quantized Onsager . Choosing the Planck constant $2\pi\hbar$ as a unit for the momentum and a space-like comoving contour we have
[…]
At the same time, the total helicity is twice the linking number of vortex loops in units of vortex circulation Moreau ; Moffat . Therefore, it is quantized in multiples of the Planck constant as an even number abanov2022axial
$\int({\bm{\pi}}\cdot\nabla\times\bm{\pi})\,d^{3}\bm{x}=2\rm Lk[vortex\ loops]\,.$ (17)
Let us write the continuity equation (4) as a particle production assisted by vorticity ‘instantons’
$\partial_{\mu}n^{\mu}=-\tfrac{k}{4}\,\Omega_{\mu\nu}{}^{\star}\Omega^{\mu\nu}\,$ (18)
[we use the relation $\partial_{\mu}h^{\mu}=\tfrac{k}{4}\,\Omega_{\mu\nu}{}^{\star}\Omega^{\mu\nu}$, where $\Omega_{\mu\nu}=\partial_{\mu}\pi_{\nu}-\partial_{\nu}\pi_{\mu}$ is the 4-vorticity tensor and ${}^{\star}\Omega^{\mu\nu}=\tfrac{1}{2}\epsilon^{\mu\nu\lambda\sigma}\Omega_{\lambda\sigma}$ is the dual tensor]. Integrating this equation over a time interval and across the entire space we find that the LHS of (18) is a change of the total particle number. The integral of the RHS is $k$ times the change of the linking number
$\Delta N=k\Delta\,\rm Lk[vortex\ loops]\,.$ (19)
Hence, an ‘instanton’ process of changing the linking of vortex loops by $1$ alters the particle number by $k$. Given that the particle number is integer, $k$ is also an integer.
Furthermore, we note that when momentum is measured in units of the Planck constant, both contributions to the current (3), should be considered as terms of comparable order in gradients.
Geometrically, Onsager’s quantization of circulation renders the gauge group (and thereby the entire fluid phase space) compact. The gauge group becomes $U(1)$ making the field $\psi$ to a phase that winds over a circle.
3. Covariant Equation of State and Natural Variables.
The existence of the covariant formulation of hydrodynamics makes derivations technically simpler and less specialized in a relativistic framework. Taking advantage of the Lorentz metric we define the particle number density $n$ through particle current as
$n^{\mu}n_{\mu}=-n^{2}\,.$ (20)
As the particle number is not conserved, the fluid exchanges particles with a reservoir, a medium of massless particles devoid of space-like momentum. We denote the density number of the reservoir constituents by $\overline{n}$, and introduce the density ratio ${S}={\overline{n}}/n$. Subsequently, the fluid energy density, being a function of $n$ should also be a function of ${S}$. This is a setting of a baroclinic flow FootnoteE .
Using (20) we treat the fluid energy density $\varepsilon(n,S)$ as a function of $n^{\mu}$. A differential of energy
$d\varepsilon=p_{\mu}dn^{\mu}+\!(\partial_{S}\varepsilon)d{S}\,.$ (21)
defines the 4-momentum $n^{\mu}$
$p_{\mu}:=\partial\varepsilon/\partial n_{\mu}=wu_{\mu}\,.$ (22)
Here $w=\partial_{n}\varepsilon$ is the analog a specific enthalpy, and $u_{\mu}:=n_{\mu}\,/n$ is a relativistic velocity, a 4-unit vector collinear to the particle current. Then we have the relation
$nw=-p_{\mu}n^{\mu}\,.$ (23)
This is a covariant version of the equation of state, which connects $p_{\mu}$ and $n^{\mu}$ FootnoteK .
Later, we will require the formula for the differential of fluid pressure, designating the fields $\{\psi,\pi,S\}$ as natural variables. In view of (23) we write the standard thermodynamic relation $-P=\varepsilon-nw$ as $-P=p_{\mu}n^{\mu}+\varepsilon$, from which obtain the formula
$-dP-\partial_{S}\varepsilon dS=n^{\mu}dp_{\mu}=$ $n^{\mu}d\pi_{\mu}-n^{\mu}\partial_{\mu}d\psi\,.$ (24)
4. Variations of Natural Variables.
Let us examine how the natural variables transform under the action of the group $\cal\ G$. The action of the gauge group is simple
$\psi\to\psi+\delta\psi\,.$ (25)
The action of the spacetime diffeomorphisms
$x^{\mu}\to x^{\mu}+\epsilon^{\mu}(x)\,$ (26)
are carried out by the Lie derivative $\cal L_{\epsilon}$ along a vector field $\epsilon$. The density ratio transform as a scalar
$\delta_{\epsilon}{S}:=\mathcal{L}_{\epsilon}{S}=\epsilon^{\mu}\partial_{\mu}{S}\,.\quad$ (27)
The momenta $\pi_{\mu}$ transform as the covector defined via a form-valued variation. Let $\pi=\pi_{\mu}dx^{\mu}$ is the momentum 1-form. Then
$\delta_{\epsilon}\pi:=\mathcal{L}_{\epsilon}\pi=\delta_{\epsilon}(\pi_{\nu}dx^{\nu})=(\delta_{\epsilon}\pi_{\nu})dx^{\nu}\,.$ (28)
Explicit form of the transformed momentum is given by the Cartan formula followed from (28)
$\delta_{\epsilon}\pi_{\nu}=\epsilon^{\mu}\partial_{\mu}\pi_{\nu}+\pi_{\mu}\partial_{\nu}\epsilon^{\mu}\,.$ (29)
5. Hamilton Principle of Hydrodynamics.
We will employ the covariant Hamiltonian principle, an efficient approach to inviscid fluid dynamics developed by Carter Carter . In this approach, the flow is characterized by the spacetime Hamilton functional $\Lambda[\pi,{S,\psi}]$, which depends on the natural variables: the gauge phase $\psi$, the canonical momentum, and the a scalar field $S$. The Hamiltonian principle asserts that on the equation of motions the Hamilton functional is invariant under the action of the group $\cal G$ Arnold . The action of the group of spacetime diffeomorphisms (27,29) can be seen as variations under physical D’Alembertian displacements of fluid parcels (26). Then the variation of the Hamilton functional vanishes on a physical flow $\delta\Lambda=0$
In this form, the Hamiltonian principle incorporates the fluid kinematics into Noether conservation laws associated with the symmetry group $\cal\ G$.
The variation of the functional
$\delta_{\epsilon}\Lambda=\int\mathcal{[J^{\mu}}\delta_{\epsilon}\pi_{\mu}+\pi_{S}\mathcal{\delta}_{\epsilon}{S}-I^{\mu}\partial_{\mu}\delta\psi]\,,$ (30)
introduces the conjugate fields: the flow current $\mathcal{J}$, a conjugate to the canonical momentum, the ‘electric’ current $I$, a conjugate to the gauge phase $\psi$, and the conjugate to the density ratio $\pi_{S}$, the time-like ‘momentum’ of reservoir’s particles. Using explicit forms of the variations (27,29) a simple algebra leads to what Carter referred to as the canonical equation of fluid motion FootnoteL
$\mathcal{J}^{\mu}\Omega_{\mu\nu}+\pi_{\nu}(\partial_{\mu}\mathcal{J}^{\mu})=\pi_{S}\,\partial_{\mu}{S}\,.$ (31)
supplemented by the conservation of the electric current (4) followed from (30) under the variation over $\psi$. The first term on the LHS is the force acting on a rotating fluid parcel. It is balanced by the force due to the fluid source (the second term sometime called the ‘ rocket term’) and the ‘heat’ source on the RHS. A notable feature of the canonical equation is that it does not make a reference to a spacetime metric.
The combined result must be gauge-invariant. The gauge phase $\psi$ introduced through the canonical momentum in the second term should not enter the equations. If the flow field $\cal J$ is gauge-invariant, its divergence vanishes and so does the second term. This is the case of the perfect fluid discussed below. However, if $\cal J$ is not gauge-invariant the $\psi$ dependence of the first term must cancel the $\psi$ dependence of the second term. This requirement imposes a nearly prohibiting condition on $\cal\ J$.
The perfect fluid is defined by the condition that the currents $\mathcal{J}$ and $I$ are both equal to $n^{\mu}$. This condition determines the Hamilton functional equal to (minus) spacetime integral of the fluid pressure Whitham ; Schutz
$\Lambda_{0}=-\int_{{\cal M}^{4}}\ P\,.$ (32)
Indeed, the differential of pressure given by (24) yields ${\cal J}^{\mu}=I^{\mu}=n^{\mu}$ and determines $\pi_{S}=\partial_{S}\varepsilon$. Then the ‘rocket term‘ is null due to the continuity condition (2) and every term in (30) is gauge-invariant. We obtain the canonical form of the Euler equation for the perfect fluid
$n^{\mu}\Omega_{\mu\lambda}=(\partial_{S}\varepsilon)\partial_{\lambda}{S}\,,$ (33)
which is equivalent to the conservation laws for the perfect fluid given by (1,2,6).
The deformation of the Hamilton functional disrupts the accidental identity of all currents and alters the mechanism that brings (30) to its gauge-invariant form.
6. Fluid Phase Space and a Generalized Hopf Fibration.
In addition to the 4-dimensional cotangent space of possible values of momentum, the phase space includes the scalar ${S}$. That makes the phase space 5-dimensional, matching the dimension of the manifold of the symmetry group $\cal G$. We illustrate this important feature by invoking the realization of the momentum through five Clebsch potentials. It suffices to consider the perfect fluid.
From the canonical equation (33) it follows that the vorticity 2-form $\Omega=\tfrac{1}{2}\Omega_{\mu\nu}({\rm{d}}x^{\mu}\wedge\!{\rm{d}}x^{\nu})$ of the baroclinic flow, where $d{S}\neq 0$, is non-degenerate ${\rm\ det}\,\Omega_{\mu\nu}\neq 0$. Hence, vorticity provides a symplectic structure to the fluid cotangent bundle. [This is in contrast to the homentropic or barotropic flow, where the RHS of (33) vanishes and $n^{\mu}$ is the null-vector of the vorticity tensor].
Then there are four local coordinates $\{\theta,e,\eta,S\}$, in which the symplectic structure (vorticity), takes on a canonical form (Darboux’s theorem): $\Omega=d\theta\wedge de+d\eta\wedge dS$. As a result, the canonical momenta are locally represented by five coordinates and that one of the coordinate is identified with $S$
$\pi=d\psi+\theta de+\eta dS\,,\,$ (34)
We are provided with a map of the 5-dimensional phase space, denoted by $\mathrm{N}^{5}$ to the 4-dimensional spacetime $\mathrm{M}^{4}$: $\mathrm{N}^{5}\to\mathrm{M}^{4}$, where a point of a spacetime $x$ is mapped out from a distinct circle $S^{1}$, represented by the gauge phase $\psi$. The local coordinates of $\mathrm{N}^{5}$ are associated with five Clebsch potentials $\{\psi,\theta,e,\eta,S\}$ Whitham ; Schutz ; Lynden-Bell . Then the physical canonical momentum $\pi=(\partial_{\mu}\psi+\theta\partial_{\mu}e+\eta\partial_{\mu}S)dx^{\mu}$, being the 1-form in $\mathrm{M}^{4}$ could be seen as a push-forward of the 1-form (34) in $\mathrm{N}^{5}$. The map describes a fibration of the phase space $S^{1}\hookrightarrow\mathrm{N}^{5}\to\mathrm{M}^{4}$. The total space ${\mathrm{N}}^{5}$ consists of fibers, with each fiber being a circle $S^{1}$ spanned by the gauge phase one for each point of the spacetime (the base of the bundle). This setup extends the classical Hopf fibration, given by the Hopf map $S^{1}\hookrightarrow S^{3}\to S^{2}$, whose extension to general manifolds is described in Novikov3 . The map is characterized by the invariant, called Hopf-Novikov invariant Khesin2 . The invariant is the integral of the top-form in $\mathrm{N}^{5}$, which is constructed from the pull-back of the canonical momentum Footnote3
$\mathrm{H}=\int_{N^{5}}\pi\wedge(d\pi)^{2},\quad(d\pi)^{2}=d\pi\wedge d\pi\,.$ (35)
It is analogous to the realization of the Hopf invariant in terms of differential forms Whitehead . In the context of semiclassical hydrodynamics the invariant is the volume of the compact phase space
$H=\tfrac{1}{2}\mathcal{H}\mathcal{N}$ (36)
[…]
7. Multivalued Functional.
The five dimensional phase space allows the following interpretation: our fluid can be seen as a boundary of an auxiliary 5d fluid. We assume that the auxiliary fluid occupies a 5-dimensional half-space $\mathrm{M}_{+}^{5}$ of a closed space $\mathrm{M}^{5}$. The boundary of the half-space is the physical spacetime ${\mathrm{M}}^{4}=\partial\mathrm{M}_{+}^{5}$. We can think about the fifth coordinate as a gauge phase $\psi$. Then the map $\mathrm{M}^{5}\to\mathrm{N}^{5}$ defines the momentum of the auxiliary fluid via (34). Would the 5d fluid occupy the entire space $\mathrm{M}^{5}$ its total helicity would be the invariant $\mathrm{H}$. The helicity of the 5d fluid occupying a half-space is the multivalued Novikov’s functional
$\Lambda_{WZN}=\int_{\mathrm{M}_{+}^{5}}\pi\wedge(d\pi)^{2}\,.$ (37)
The integrand in (37) is a Jacobian of map $\mathrm{M}^{5}\to\mathrm{N}^{5}$, a closed form $\pi\wedge(d\pi)^{2}=d\Phi$ making the integral a surface term spanned over physical spacetime
$\Lambda_{WZN}=-\int_{{\mathrm{M}}^{4}}\Phi\quad({\rm mod}\,\mathrm{H})\,$ (38)
modulo the invariant (35) [we set $\Phi$ to vanish at a reference point in the bulk of $\mathrm{M}_{+}^{5}$ and run the integralin (37) from the boundary to this point]. In this sense the functional is multivalued. Consequently, $\Phi$ can not be expressed in a coordinate-free manner. Still, it could be elementary computed in chosen coordinates. Choosing the gauge phase as a fifth coordinate $(\psi,\,x^{\mu})$, the density $\Phi$, modulo an exact 4-form, is
$\Phi\approx\psi(d\pi)^{2}\approx-\partial_{\mu}\psi\,h^{\mu}\,\approx-\partial_{\mu}\psi\,\Sigma^{\mu}.$ (39)
8. Multivalued Functional in Fluid dynamics.
We employ the multivalued functional to deform the Hamilton functional of the perfect fluid
$\Lambda=\Lambda_{0}+\tfrac{k}{2}\Lambda_{WZN}=-\int_{M^{4}}[P+\tfrac{k}{2}\Phi]\,.$ (40)
While the added functional is not uniquely defined, it nonetheless generates local equation of motion (cf.,Novikov2 ). The ambiguity of the functional does not extend to its variation as the invariant $\mathrm{H}$ does not vary.
Unlike $\Lambda_{0}$, the functional is not gauge-invariant. It opens a channel of inflow of the 5d auxiliary fluid into the physical fluid. Nevertheless, the equations of motion maintain gauge invariance.
In the semiclassical fluid, the multivaluedness of $\Lambda$ leads to the quantization of $k$, as we already discussed in the introduction. It follows from the requirement for $e^{(i/\hbar)\Lambda}$ to be single-valued under a global gauge transformation. Consider a transformation $\psi\to\psi+1$, applied at a fixed time. It changes the functional by the total helicity $\Lambda\to\Lambda+\tfrac{k}{2}\int({\bm{p}}\cdot\nabla\times\bm{p})\,d^{3}\bm{x}$. Since the latter is an even integer, $k$ is quantized (cf.,Witten ).
9. Euler Equation.
Now we turn to equations of motions generated by the deformed Hamilton functional. We calculate the Noether currents as defined by (30), and subsequently, substitute them into the continuity equation (3) and the canonical equation (31). First, we vary over $\psi$ while holding $\pi$ fixed. With the help of (39) and (24) we obtain the formula for the ‘electric’ current (3) and deformed continuity equation (18).
The next step is to consider a variation of (40) over the canonical momentum. This yields ${\cal\ J^{\mu}}=n^{\mu}+k\epsilon^{\mu\nu\lambda\sigma}\partial_{\nu}\psi\,\partial_{\lambda}\pi_{\sigma}$. The flow field is not gauge-invariant, however, its divergence $\partial_{\mu}\mathcal{J}^{\mu}=\partial_{\mu}n^{\mu}$ is. Now we have all the components to specialize the canonical equation (31). The remaining step is to verify that the gauge phase $\psi$ cancels out: We can see it with the help of the identity
$\epsilon^{\mu\nu\lambda\sigma}z_{\rho}+\cdots 0\,,$ (41)
where the ellipsis denotes the cyclic permutation of all five indices, and $z_{\mu}$ is an arbitrary 5-covector. The result is the gauge-invariant canonical form of the Euler equation
$n^{\mu}\Omega_{\mu\nu}+p_{\nu}(\partial_{\mu}n^{\mu})=(\partial_{S}\varepsilon)\,\partial_{\mu}{S}\,$ (42)
which is equivalently represented by the Newtonian form:
$n^{\mu}\partial_{\mu}p_{\nu}+\partial_{\nu}P=-p_{\nu}(\partial_{\mu}n^{\mu})\,,$ (43)
or by the energy-momentum conservation laws (1,6).
10. Spin, Spin Current and Spin-Orbit Coupling.
With the help of the continuity equation (18) we reformulate the RHS of (43) as a force
$n^{\mu}\partial_{\mu}p_{\nu}+\partial_{\nu}P=-k\Sigma^{\mu}\Omega_{\mu\nu}\,.$ (44)
Now, it becomes apparent that our fluid is spinning generating a totally antisymmetric spin current $\Sigma_{\mu\nu\lambda}=\frac{1}{2}\epsilon_{\mu\nu\lambda\sigma}\Sigma^{\sigma}$ with helicity being a spin. Given that the vorticity is twice the angular velocity of a rotating fluid parcel, the force in the RHS of (44) represents the spin-orbit coupling. The WZN term gives fluid particles a spin.
In a separate publication, we demonstrate that our hydrodynamics aligns with the kinematics of Weyl fermions, where the integer parameter $k$ is the number of fermionic species, and $\Sigma^{\mu}$ is the fermionic spin.
11. Advection.
With the help of the continuity equation (18) and the identity (41) we write Eq.(42) as
$(n^{\mu}-k\Sigma^{\mu})\Omega_{\mu\nu}=(\partial_{S}\varepsilon)\,\partial_{\mu}{S}\,.$ (45)
The form of the Euler equation equation is convenient to obtain advection property of the Clebsch potentials. As expected they are the same as in the perfect fluid
$\mathcal{L}\theta=\mathcal{L}e=\mathcal{L}S=0,$ (46)
$\mathcal{L}\eta=\partial_{S}\varepsilon,\quad\mathcal{L}\psi=-nw\,,$ (47)
except that now the Lie derivative $\mathcal{L}=(n^{\mu}-k\Sigma^{\mu})\partial_{\mu}$ acts along the deformed vector field $n^{\mu}-k\Sigma^{\mu}$. Contrary to the perfect fluid this field is not conserved.
12. Homentropic Flows.
A homentropic flow occurs when the density ratio $S$ is uniform and constant. In this situation, and also in barotropic flows the vorticity tensor is degenerate ${\rm\ det}\,\Omega_{\mu\nu}=0$, having rank 2. This prevents the construction of the WZN-term since the phase space of a homentropic flow is not symplectic. The homentropic flow conserves helicity $\Sigma^{\mu}$, and independently conserves the particle number $n^{\mu}$ providing no exchange with the reservoir. In this case, the equations motion are no different from that of the perfect fluid. This can be observed from the relation $\Sigma^{\lambda}\partial_{\lambda}S=\tfrac{1}{2}\left({d\bar{n}}/{dn}\right)_{\varepsilon}\,\partial_{\mu}\Sigma^{\mu}$ which follows from (45) [the derivative is taken at a fixed energy]. When $S$ is constant, both $\partial_{\mu}\Sigma^{\mu}$ and the WZN term vanish.
The work was supported by the NSF under Grant NSF DMR-1949963. He acknowledges insightful discussions with V. P. Nair, G. Volovik, M. Stephanov, B. Shapiro, L. Friedlander, and is especially grateful to A. G. Abanov and A. Cappelli for collaboration on this subject.